# The generalized Ramsey number $f(n,5,8)=\frac{6}{7}n+o(n)$

### v1

@@ 5.2. Checking conditions of Theorem 2
Much of the content of Lemma 1 is heuristically easy to see (and much of it is easy to prove). The proof of lemma 1 is in Appendix A, and here we will just explain why it is believable. We already know Phase A colors almost all edges so part (1) makes sense. About $n^{2}/2$ edges are colored and about $n^{2-\delta}/2$ are sharable and so part (2) makes sense. To heuristically see part (3), first note that we need to choose $a-2$ vertices to complete a set $S\in\mathcal{S}_{a,b}(uv)$. Then recall that the Phase A coloring is the result of a random process in which colored gadgets are chosen one at a time, and heuristically the probability that one randomly chosen color matches another one (i.e. the probability of each color repeat) is on the order $1/n$. Part (4) is, as we will see in Appendix A, bounding a family of configurations of gadgets each of which is heuristically (and actually) unlikely to appear a lot.
@@ 6. Phase B
Let $C_{B}:=C_{B1}\cup C_{B2}$ where $C_{B1},C_{B2}$ are two disjoint sets of set of $2d_{B}$ new colors each. $Y$ will contain a vertex $e_{c}$ for every $(e,c)\in E^{\prime\prime}\times C_{B1}$ and for every $(e,c)\in E^{\prime\prime\prime}\times C_{B2}$. Condition (H1B) is satisfied since $\Delta_{2}(\mathcal{H}_{B})=1$. Condition (H2B) is satisfied since every vertex in $X$ has degree $2d_{B}$ and every vertex in $Y$ has degree 1. An $X$-perfect matching in $\mathcal{H}_{B}$ corresponds to a coloring of the uncolored edges in the obvious way. Note that the sharable edges (in $E^{\prime\prime}$) only get colors from $C_{B1}$ and the nonsharable edges (in $E^{\prime\prime\prime}$) get colors from $C_{B2}$.
@@
Let $\mathcal{S}_{a,b}$ be the family of sets $S\subseteq V(K_{n})$ with $|S|=a$ and having $b$ color repetitions in the coloring from Phase A. To define $\mathcal{C}_{B}$ we will first define another conflict system $\mathcal{D}$ and then let $\mathcal{C}_{B}$ be the set of minimal (inclusion-wise) conflicts of $\mathcal{D}$. $\mathcal{D}$ will have the following conflicts, described in order of increasing size.
@@
  • $f$ is in some $\mathcal{S}_{4,1}(e)$, or
  • $f$ is in some $\mathcal{S}_{5,2}(e)$.
@@
  • $f,e^{\prime},f^{\prime}$ are in some $\mathcal{S}_{4,0}(e)$, or
  • $f,e^{\prime},f^{\prime}$ are in some $\mathcal{S}_{5,1}(e)$.
6-conflicts in $\mathcal{D}$: $\{e_{c},f_{c},e^{\prime}_{c^{\prime}},f^{\prime}_{c^{\prime}},e^{\prime\prime}_{c^{\prime\prime}},f^{\prime\prime}_{c^{\prime\prime}}\}$ will be a 6-conflict if $f,e^{\prime},f^{\prime},e^{\prime\prime},f^{\prime\prime}$ are all in some $\mathcal{S}_{5,0}(e)$.
@@
We check condition (C1B), starting with $j=2$. Fix $e_{c}$ and we will bound the number of $f_{c}$ such that $\{e_{c},f_{c}\}$ is a conflict. There are $O(n^{1-\varepsilon_{A}^{3}})$ edges $f$ adjacent to $e$ by part (1), and $O(n)$ edges $f$ in some $\mathcal{S}_{4,1}$ or $\mathcal{S}_{5,2}$ by part (3). Thus we have $\Delta(\mathcal{D}^{(2)})=O(n)=o(d_{B}\log d_{B})$. Now we check $j=4$. We have $\Delta(\mathcal{D}^{(4)})=O(n^{3})=o(d_{B}^{3}\log d_{B})$ since, fixing some $e_{c}$, we can determine the rest of a 4-conflict by first choosing some $\mathcal{S}_{4,0}$ or $\mathcal{S}_{5,1}$ which are $O(n^{2})$ by part (3) and then choosing a color $c^{\prime}$. For $j=6$ using part (3) we similarly have $\Delta(\mathcal{D}^{(6)})=O(n^{5})=o(d_{B}^{5}\log d_{B})$. Thus (C1B) holds for any positive $\alpha$.
@@
Indeed, consider the case of a 4-conflict in $\mathcal{C}_{B}$ arising from some $\mathcal{S}_{4,0}(e)$. If we fix $j^{\prime}=2$ of these colored edges then we have fixed either $2$ colors and $3$ vertices or $1$ color and $4$ vertices, and in either case there are $O(n)$ ways to complete the alternating 4-cycle. If we fix $j^{\prime}=3$ colored edges they are in at most one alternating 4-cycle. Now consider the case of a 4-conflict in $\mathcal{C}_{B}$ arising from some $\mathcal{S}_{5,1}(e)$. We can assume that this conflict does not make an alternating 4-cycle since we just handled that case (then we would have an $\mathcal{S}_{4,0}(e)$). Thus the 4 colored edges in $K_{n}$ corresponding this conflict form an alternating 4-path. If we fix $j^{\prime}=2$ of these colored edges there are $O(n)$ ways to complete the conflict. If we fix $j^{\prime}=3$ of them then by Lemma 1 part (1) there are $O(n^{1-\varepsilon_{A}^{3}})$ ways to complete the conflict. Thus we have checked (C2B) for $j=4$ (this condition just has to hold for some $\varepsilon=\varepsilon_{B}>0$ and we can choose $\varepsilon_{B}\ll\varepsilon_{A},\delta$).
@@
Now we check (C3B). Let $v\in V(\mathcal{H}_{B}),e\in E(\mathcal{H}_{B})$ with $v\notin e$. Recall that $e$ corresponds to coloring some color $c_{e}$ to some edge $f\in E(K_{n})$. It is only possible to have a 2-conflict $\{e,e^{\prime}\}\in\mathcal{D}$ if $e^{\prime}$ corresponds to assigning some edge $f^{\prime}\in E(K_{n})$ the same color $c_{e}$. Thus,
|$\big{|}\big{\{}e^{\prime}\in E(\mathcal{H}_{B}):e^{\prime}\ni v\mbox{ and }\{e,e^{\prime}\}\in\mathcal{D}\big{\}}\big{|}\leq 1$
and (C3B) holds.
Now we check (C4B). Fix $e,e^{\prime}\in E(\mathcal{H}_{B})$. By Lemma 1 part (4),
|$\big{|}\big{\{}e^{\prime\prime}\in E(\mathcal{H}_{B}):\{e,e^{\prime\prime}\}\in\mathcal{D}\mbox{ and }\{e^{\prime},e^{\prime\prime}\}\in\mathcal{D}\big{\}}\big{|}\leq n^{1-\delta/2}.$
@@ A.4. Proof of Lemma 1(4)
We consider cases based on what role $e^{\prime}$ plays with $e_{1}$ and with $e_{2}$. We have $e^{\prime}\in\mathcal{S}_{4,1}(e_{1})\cup\mathcal{S}_{5,2}(e_{1})$. Figure 10(3) shows the case $e^{\prime}\in\mathcal{S}_{4,1}(e_{1})$, and 10(7)-(16) shows all possible cases $e^{\prime}\in\mathcal{S}_{5,2}(e_{1})$ (where $e^{\prime}$ is one of the dashed edges and $e_{1}=uv$). Likewise, $e^{\prime}$ also plays one of these roles with respect to $e_{2}$. Thus, the cases we consider are all possible combinations of two roles (one for $e_{1}$, one for $e_{2}$) $e^{\prime}$ could play. The role $e^{\prime}$ plays with each of $e_{1},e_{2}$ implies that we certain colored edges which must come from some set of gadgets. For each case, the family of these sets of gadgets (which together would cause $e^{\prime}$ to play the prescribed role with $e_{1}$ and $e_{2}$) will be one of our hypergraphs $\mathcal{X}$ when we apply Lemma 2.

### v2

@@ 5.2. Checking conditions of Theorem 2
Much of the content of Lemma 1 is heuristically easy to see (and much of it is easy to prove). The proof of lemma 1 is in Appendix A, and here we will just explain why it is believable. We already know Phase A colors almost all edges so part (1) makes sense. About $n^{2}/2$ edges are colored and about $n^{2-\delta}/2$ are sharable and so part (2) makes sense. To heuristically see part (3), first note that to extend $uv$ to a set $S\in\mathcal{S}_{a,b}(uv)$ we need to choose $a-2$ vertices. Then recall that the Phase A coloring is the result of a random process in which colored gadgets are chosen one at a time, and heuristically the probability that one randomly chosen color matches another one (i.e. the probability of each color repeat) is on the order $1/n$. Part (4) is, as we will see in Appendix A, bounding a family of configurations of gadgets each of which is heuristically (and actually) unlikely to appear a lot.
@@ 6. Phase B
Let $C_{B}:=C_{B1}\cup C_{B2}$ where $C_{B1},C_{B2}$ are two disjoint sets of $2d_{B}$ new colors each. $Y$ will contain a vertex $e_{c}$ for every $(e,c)\in E^{\prime\prime}\times C_{B1}$ and for every $(e,c)\in E^{\prime\prime\prime}\times C_{B2}$. Each $e\in X=E^{\prime\prime}\cup E^{\prime\prime\prime}$ will be adjacent to every $e_{c}\in Y$. For clarity, $\mathcal{H}_{B}$ is just the disjoint union of stars, where each $e$ is the center of a star which has the $e_{c}$ as leaves. For $e\in X$ and $e_{c}\in Y$, we abuse notation by writing $e_{c}$ to also denote the edge $ee_{c}$ of $\mathcal{H}_{B}$. Condition (H1B) is satisfied since $\Delta_{2}(\mathcal{H}_{B})=1$. Condition (H2B) is satisfied since every vertex in $X$ has degree $2d_{B}$ and every vertex in $Y$ has degree 1. An $X$-perfect matching in $\mathcal{H}_{B}$ corresponds to a coloring of the uncolored edges in the obvious way. Note that the sharable edges (in $E^{\prime\prime}$) only get colors from $C_{B1}$ and the nonsharable edges (in $E^{\prime\prime\prime}$) get colors from $C_{B2}$.
@@
To define $\mathcal{C}_{B}$ we will first define another conflict system $\mathcal{D}$ and then let $\mathcal{C}_{B}$ be the set of minimal (inclusion-wise) conflicts of $\mathcal{D}$. $\mathcal{D}$ will have the following conflicts, described in order of increasing size.
@@
  • $f$ is in some $S\in\mathcal{S}_{4,1}(e)$ (see Figure 10(1)), or
  • $f$ is in some $S\in\mathcal{S}_{5,2}(e)$.
@@
  • $f,e^{\prime},f^{\prime}$ are in some $S\in\mathcal{S}_{4,0}(e)$ (see Figure 10(2)), or
  • $f,e^{\prime},f^{\prime}$ are in some $S\in\mathcal{S}_{5,1}(e)$ (see Figure 10(3)).
6-conflicts in $\mathcal{D}$: $\{e_{c},f_{c},e^{\prime}_{c^{\prime}},f^{\prime}_{c^{\prime}},e^{\prime\prime}_{c^{\prime\prime}},f^{\prime\prime}_{c^{\prime\prime}}\}$ will be a 6-conflict if $f,e^{\prime},f^{\prime},e^{\prime\prime},f^{\prime\prime}$ are all in some $S\in\mathcal{S}_{5,0}(e)$.
$e^{\prime}$$e$
1
$e^{\prime}$$f$$e$$f^{\prime}$
2
$e$$e^{\prime}$$f$$f^{\prime}$
3
Figure 10. Representations of a 2-conflict and two 4-conflicts in $\mathcal{D}$. All colored edges are from Phase A, and dashed edges will be colored in Phase B.
@@
Proposition 2.
A $\mathcal{C}_{B}$-free $X$-perfect matching in $\mathcal{H}_{B}$ corresponds to a $(5,8)$-coloring of $K_{n}$.
Proof.
Recall that Phase A produces a partial $(5,8)$-coloring, so in particular the Phase A coloring has at most 2 color repetitions in any set of 5 vertices. Now suppose we have a $\mathcal{C}_{B}$-free $X$-perfect matching in $\mathcal{H}_{B}$. Since our matching is $\mathcal{C}_{B}$-free it is also $\mathcal{D}$-free. Since $X=E^{\prime\prime}\cup E^{\prime\prime\prime}$, our matching colors every edge that did not get colored in Phase A. To show that the final coloring is a $(5,8)$-coloring, let $S$ be a set of 5 vertices. From Phase A, $S$ has either 0, 1, or 2 repeats. If $S\notin\mathcal{S}_{5,0}(e)\cup\mathcal{S}_{5,1}(e)\cup\mathcal{S}_{5,2}(e)$ for all $e$, this can only be because $S$ does not contain two disjoint edges that are both in $E^{\prime\prime}$ or both in $E^{\prime\prime\prime}$ (see Definition 4). In that case it is not possible for $S$ to gain any new color repetitions in Phase B. If $S\in\mathcal{S}_{5,0}(e)$ for some $e$ then $S$ cannot get 3 new repetitions in Phase B since that situation would arise from a 6-conflict (note that in a $\mathcal{D}$-free matching we cannot have 3 edges of the same Phase B color in $S$ since our 2-conflicts enforce that adjacent edges have different colors). If $S\in\mathcal{S}_{5,1}(e)$ for some $e$ then $S$ cannot get 2 new repetitions in Phase B since that would be a 4-conflict. Finally, if $S\in\mathcal{S}_{5,2}(e)$ for some $e$ then $S$ cannot get any new repetitions in Phase $B$ since that would be a 2-conflict. ∎
We check condition (C1B), starting with $j=2$. Fix $e_{c}$ and we will bound the number of $f_{c}$ such that $\{e_{c},f_{c}\}$ is a conflict. There are $O(n^{1-\varepsilon_{A}^{3}})$ edges $f$ adjacent to $e$ by part (1), and $O(n)$ edges $f$ in some $S\in\mathcal{S}_{4,1}(e)\cup\mathcal{S}_{5,2}(e)$ by part (3). Thus we have $\Delta(\mathcal{D}^{(2)})=O(n)=o(d_{B}\log d_{B})$. Now we check $j=4$. We have $\Delta(\mathcal{D}^{(4)})=O(n^{3})=o(d_{B}^{3}\log d_{B})$ since, fixing some $e_{c}$, we can determine the rest of a 4-conflict by first choosing some $S\in\mathcal{S}_{4,0}(e)\cup\mathcal{S}_{5,1}(e)$ which are $O(n^{2})$ by part (3) and then choosing a color $c^{\prime}$. For $j=6$ using part (3) we similarly have $\Delta(\mathcal{D}^{(6)})=O(n^{5})=o(d_{B}^{5}\log d_{B})$. Thus (C1B) holds for any positive $\alpha$.
@@
Indeed, consider the case of a 4-conflict in $\mathcal{C}_{B}$ arising from some $S\in\mathcal{S}_{4,0}(e)$. If we fix $j^{\prime}=2$ of these colored edges then we have fixed either $2$ colors and $3$ vertices or $1$ color and $4$ vertices, and in either case there are $O(n)$ ways to complete the alternating 4-cycle. If we fix $j^{\prime}=3$ colored edges they are in at most one alternating 4-cycle. Now consider the case of a 4-conflict in $\mathcal{C}_{B}$ arising from some $S\in\mathcal{S}_{5,1}(e)$. We can assume that this conflict does not make an alternating 4-cycle since we just handled that case (then we would have $S\in\mathcal{S}_{4,0}(e)$). Thus the 4 colored edges in $K_{n}$ corresponding this conflict form an alternating 4-path. If we fix $j^{\prime}=2$ of these colored edges there are $O(n)$ ways to complete the conflict. If we fix $j^{\prime}=3$ of them then by Lemma 1 part (1) there are $O(n^{1-\varepsilon_{A}^{3}})$ ways to complete the conflict. Thus we have checked (C2B) for $j=4$ (this condition just has to hold for some $\varepsilon=\varepsilon_{B}>0$ and we can choose $\varepsilon_{B}\ll\varepsilon_{A},\delta$).
@@
Now we check (C3B). We fix a vertex and a non-incident edge in $\mathcal{H}_{B}$. If our vertex is in $Y$ then it is in at most one edge of $\mathcal{H}_{B}$ and we are done. So assume our fixed vertex is say $e\in X$, and our fixed edge is say $f_{c}$. The only edge containing $e$ in $\mathcal{H}_{B}$ that could form a conflict with $f_{c}$ is $e_{c}$ since 2-conflicts only have one color. Thus (C3B) holds.
Now we check (C4B). Fix $e_{c},e^{\prime}_{c}\in E(\mathcal{H}_{B})$. By Lemma 1 part (4),
|$\big{|}\big{\{}e^{\prime\prime}_{c}\in E(\mathcal{H}_{B}):\{e_{c},e^{\prime\prime}_{c}\}\in\mathcal{D}\mbox{ and }\{e^{\prime}_{c},e^{\prime\prime}_{c}\}\in\mathcal{D}\big{\}}\big{|}\leq n^{1-\delta/2}.$
@@ A.4. Proof of Lemma 1(4)
We consider cases based on what role $e^{\prime}$ plays with $e_{1}$ and with $e_{2}$. We have $e^{\prime}\in\mathcal{S}_{4,1}(e_{1})\cup\mathcal{S}_{5,2}(e_{1})$. Figure 11(3) shows the case $e^{\prime}\in\mathcal{S}_{4,1}(e_{1})$, and 11(7)-(16) shows all possible cases $e^{\prime}\in\mathcal{S}_{5,2}(e_{1})$ (where $e^{\prime}$ is one of the dashed edges and $e_{1}=uv$). Likewise, $e^{\prime}$ also plays one of these roles with respect to $e_{2}$. Thus, the cases we consider are all possible combinations of two roles (one for $e_{1}$, one for $e_{2}$) $e^{\prime}$ could play. The role $e^{\prime}$ plays with each of $e_{1},e_{2}$ implies that we certain colored edges which must come from some set of gadgets. For each case, the family of these sets of gadgets (which together would cause $e^{\prime}$ to play the prescribed role with $e_{1}$ and $e_{2}$) will be one of our hypergraphs $\mathcal{X}$ when we apply Lemma 2.